A small block slides without friction down an inclined starting form rest. Let Sn be the distance travelled from time


t = n-1 to t = n. Then Sn/Sn+1 is -


(A)   2n-1/2n


(B)   2n+1/2n+1


(C)   2n-1/2n+1


(D)   2n/2n+1

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear aditya,

The correct answer is c)

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

Sukhendra Reddy Rompally B.Tech Mining Machinery Engg, ISM Dhanbad
13 years ago
Hi Aditya! The distance travelled by a body during the nth second of its motion is given by Sn=u+a(n-0.5),where u is the initial velocity of the particle,and a is the uniform acceleration.This equation can be derived quite easily by calculating the distance travelled in first n sec and first n-1 sec and subtracting them.Here u=0,a=gsinx,where x is the angle of inclination.now substituting the values,we get the required ratio=(n-0.5)/(n+1 -0.5) = (2n-1)/(2n+1) ALL THE BEST PLZZ APPROVE MY ANSWER IF YOU LIKE IT